OpenStudy (anonymous):

What is the sum of the arithmetic sequence 174, 168, 162 …, if there are 35 terms?

OpenStudy (anonymous):

Actually .. I need more help with this one? A monument is built out of toothpicks. There are 25 rows and each row has eight more toothpicks than the row above. The top row has five toothpicks. How many toothpicks are on the bottom row? Using complete sentences, explain the procedure taken to answer this question.

OpenStudy (anonymous):

Do you understand ven?

OpenStudy (anonymous):

a1=174 a2=168 d=a2-a1 d=168-174=-6 Sn=n/2(2a+(n-1)d) S35=35/2(2*174+34(-6) =17.5(348-204) =2520

OpenStudy (anonymous):

@Serena:Looking at those.

OpenStudy (anonymous):

So, there are 25 rows Last row=l=5 toothpicks... You know there's a formula somewhere.. I don't remember it though :( sorry..